HUBBLE SPACE TELESCOPE DAILY REPORT #2098 PERIOD COVERED: 0000Z (UTC) 04/08/98 - 0000Z (UTC) 04/09/98 Daily Status Report as of 099/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.2 Completed Two Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Nine Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.4 Completed STIS/CCD 7661 (STIS MAMA Filter Red Leak Measurement) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of standard star BD75D325 to make filter red leak calibration measurements. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed NIC/3 7499 (NICMOS Extragalactic Parallel Grism Survey) The Near Infrared Camera (Camera Three) was used to make observations as part of an extragalactic parallel Grism Survey. This proposal is a parallel program using NICMOS's unique capabilities for slitless spectroscopy in the near IR. The observations were completed as planned, and no anomalies were reported. 1.6 Completed STIS/CCD 7803 (Daily Darks to Update Acquisition Bad Pixel Table Part III) The Space Telescope Imaging Spectrograph (CCD) was used to make daily dark calibration updates. These daily darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were completed as planned, and no anomalies were reported. 1.7 Completed Three Sets of WF/PC-2 7336 (Cosmological Parameters Omega and Lambda From High-Redshift Type Ia Supernovae) The WF/PC-2 was used to make observations of the high redshifted Type Ia supernovae 14H0.5Z2, 11H0.5Z1, 11H0.5Z2 and 14H0.5Z1. HST can make a direct measurement of the cosmological parameters, Omega, Lambda, and curvature, using Type Ia supernovae as calibrated "standard candles." The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed Seven Sets of NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-= 98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. The observations were completed as planned, and no anomalies were reported. 1.9 Completed STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain =3D 1 and gain =3D 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed Two Sets of WF/PC-2 7590 (Cosmological Parameters Omega and Lambda From High-Redshift Type IA Supernovae) The WF/PC-2 was used to make observations of type 1A supernova 11H0.9Z1 and 11H0.9Z2 in order to study their redshifts. The HST was used to make direct measurements of the cosmological parameters Omega, Lambda, and curvature using Type 1A supernovae as a calibrated "standard candle". The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed NIC/2 7330 (The Fueling of Active Nucleii: A NICMOS Snapshot Survey) The Near Infrared Camera (Camera Two) was used to probe the dark matter distribution of NGC-3300 far from the center by observing intervening absorption from the MgII doublet (lambda 2800) in the spectrum of one background QSO located in a geometrically optimal position. HST observations are required because this feature is in a wavelength range not accessible from the ground, and the QSO magnitudes and expected strength of absorption make IUE observations unfeasible. The proposal completed as expected with no anomalies. 1.12 Completed NIC/1/2/3 7423 (CRL2688: The Rosetta Stone for the Evolution of AGB Red Giants into Bipolar Planetary Nebulae) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of CLR2688. This object is considered to be the Rosetta Stone for the evolution of AGB red giants into bipolar planetary nebulae. This detailed study promises to lead to major breakthroughs in our understanding of late stellar evolution. The observations were completed as planned, and no problems were reported. 1.13 Completed STIS/CCD 7911 (STIS Galactic Plane Pure Parallel Program) The Space Telescope Imaging Spectrograph (CCD) was used to make low galactic latitude survey observations in parallel with the prime science observations. Short exposures with the 50CCD imaging mode will provide the wavelength zero-point calibration. These data will be placed immediately into the Hubble Data Archive. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 11 Successful : 11 Scheduled Reacquisitions : 3 Successful : 3 2.2 FHST Updates: Scheduled : 27 Successful : 27 2.3 Operations Notes: At 098/1714Z, the MA2 transmitter was turned on for a "blind acquisition" in accordance with ROP IC-01. The STIS MAMA-1 MCE reset at 098/184950Z while in low voltage and while outside an SAA contour. HSTAR 6432 was written. The STIS flighjt software cpunter was reset using ROP NS-12 a=DD 098/2006Z. Recovery will be accomplished via the SMS at the time of the next high voltage on event. Per ROP SR-01, the SSR EDAC errors were cleared on two occasions. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH